OpenStudy (anonymous):

if one zero and poly.3x^-8x-2k-1 is seven times the other.Find the zeroes and value of k.

OpenStudy (anonymous):

Is it \[3x ^{-8x-2k-1}\]

OpenStudy (anonymous):

no 3x^-8x-2k-1

OpenStudy (anonymous):

3x^? nothing

OpenStudy (lgbasallote):

let's try again... \[3x^{-8x} -2k -1?\]

OpenStudy (anonymous):

is it what @lgbasallote said

OpenStudy (anonymous):

3x square . ^ means aquare

OpenStudy (anonymous):

nope 3x^2 = 3 square

OpenStudy (anonymous):

if one zero and poly.3x^-8x-2k-1 is seven times the other meaning

OpenStudy (anonymous):

@apoorvk can u do dis question fully ma frnd????

OpenStudy (anonymous):

suppose one root is a , then the other is 7a first put a in the equation 3a^2-8a-2k-1=0 then put 7a in the equation 3(7a)^2-56a-2k-1=0 then solve it

OpenStudy (anonymous):

The roots are 7 and 1 and the value of k is -3.

OpenStudy (anonymous):

hw u gt can u jst show it @adarsh\